# Training-Free Text-Guided Color Editing with Multi-Modal Diffusion Transformer

### v1

@@ Abstract.
Text-guided color editing in images and videos is a fundamental yet unsolved problem, requiring fine-grained manipulation of color attributes, including albedo, light source color, and ambient lighting, while preserving physical consistency in geometry, material properties, and light-matter interactions. Existing training-free methods offer broad applicability across editing tasks but struggle with precise color control and often introduce visual inconsistency in both edited and non-edited regions. In this work, we present ColorCtrl, a training-free color editing method that leverages the attention mechanisms of modern Multi-Modal Diffusion Transformers ($M$-DiT). By disentangling structure and color through targeted manipulation of attention maps and value tokens, our method enables accurate and consistent color editing, along with word-level control of attribute intensity. Our method modifies only the intended regions specified by the prompt, leaving unrelated areas untouched. Extensive experiments on both SD3 and FLUX.1-dev demonstrate that ColorCtrl outperforms existing training-free approaches and achieves state-of-the-art performances in both edit quality and consistency. Furthermore, our method surpasses strong commercial models such as FLUX.1 Kontext Max and GPT-4o Image Generation in terms of consistency. When extended to video models like CogVideoX, our approach exhibits greater advantages, particularly in maintaining temporal coherence and editing stability. Finally, our method also generalizes to instruction-based editing diffusion models such as Step1X-Edit and FLUX.1 Kontext dev, further demonstrating its versatility. Here is the project website.
copyright: nonecopyright: nonecopyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOGisbn: 978-1-4503-XXXX-X/2018/06submissionid: 1683
@@ 1. Introduction
The recent architectural shift in diffusion models from U-Net (Rombach et al., 2022) to the Multi-Modal Diffusion Transformer ($M$-DiT) (Esser et al., 2024) opens up new opportunities for training-free color editing, for two main reasons: (1) the new architecture allows for scaling up both data and model capacity, enabling better adherence to physical priors; and (2) the improved fusion of text and vision modalities supports more flexible and precise attention-control editing strategies.
In this work, we propose ColorCtrl, a training-free, open-world color editing method that effectively leverages pre-trained $M$-DiT models to perform natural and precise color modifications, while preserving all other visual attributes, such as geometry, material, reflection and refraction, light source position, illumination direction, and light intensity, as shown in Fig. 1. For example, when editing the color of a ball, ColorCtrl accurately adjusts not only the ball itself but also its reflection in the water, the specular reflections on both sides, and even the small water droplets on the glass surface, as demonstrated in the fourth example of Fig. 1. Furthermore, ColorCtrl supports fine-grained, word-level control over the strength of color attributes while preserving the other visual attributes mentioned above. Thanks to our precise and robust control, ColorCtrl requires no manual parameter tuning and can be directly applied across all attention layers and inference timesteps.
@@
  • We present extensive experiments demonstrating that ColorCtrl achieves state-of-the-art performance among training-free methods on $M$-DiT-based models. Compared to commercial models (i.e., FLUX.1 Kontext Max and GPT-4o Image Generation), our method delivers significantly better consistency preservation and produces more natural edits.
  • ColorCtrl generalizes well across multiple $M$-DiT-based models, including video and instruction-based editing models, highlighting its broad applicability and extensibility.
@@ 2. Related Work
Text-to-image and video generation. Diffusion models with a U-Net backbone (Ho et al., 2020; Rombach et al., 2022; Guo et al., 2024) have largely replaced early GAN systems (Reed et al., 2016; Yu et al., 2023; Wang et al., 2023) due to superior image fidelity. However, the U-Net design scales poorly, leading to the adoption of Diffusion Transformers (DiT) (Peebles and Xie, 2023). Among these, $M$-DiT (Esser et al., 2024) has emerged as a widely adopted backbone for recent state-of-the-art models (Esser et al., 2024; AI, 2024; Labs, 2024; Yang et al., 2024; Kong et al., 2024; Liu et al., 2025b; Black Forest Labs, 2025), such as SD3 (Esser et al., 2024) and FLUX.1-dev (Labs, 2024) for image generation, as well as CogVideoX (Yang et al., 2024) for video generation. In this work, we propose an attention-control scheme that plugs into any $M$-DiT model.
Text-guided editing. Training-free text-guided editing methods based on pre-trained diffusion models offer flexibility and efficiency. Existing approaches fall into two groups: (i) sampling-based methods, which steer generation by injecting controlled noise or inversion (Jiao et al., 2025; Huberman-Spiegelglas et al., 2024; Xu et al., 2023; Kulikov et al., 2024; Yan et al., 2025); (ii) attention-based methods, which modify attention maps, starting with Prompt-to-Prompt (Hertz et al., 2023) and its image/video variants (Chen et al., 2024; Wang et al., 2025; Liu et al., 2024; Cao et al., 2023; Cai et al., 2025; Rout et al., 2025; Xu et al., 2025). Among these, DiTCtrl is the only method exploring attention control in $M$-DiT, but ColorCtrl differs in key ways: DiTCtrl only applies mask extraction during long video generation, not color editing, while our improved method is more robust. Additionally, DiTCtrl’s attribute re-weighting disrupts attention consistency, causing inconsistent geometries, while our method avoids this. Furthermore, ColorCtrl works across all layers and timesteps, unlike DiTCtrl, which requires careful layer selection. Different from DiTCtrl, we operate in attention maps rather than key and value tokens. More recently, methods (Liu et al., 2025a; Brooks et al., 2023) like FLUX.1 Kontext (Black Forest Labs, 2025) and GPT-4o (OpenAI, 2025) have shown convenient creative workflows in training with synthetic instruction-response pairs for fine-tuning a diffusion model for image editing.
@@ 3. Method
We first formulate the color editing problem in Sec. 3.1, followed by a revisit of the attention mechanism in $M$-DiT blocks in Sec. 3.2. Sec. 3.3 describes our approach for preserving geometry, material properties, and light-matter interactions. In Sec. 3.4, we introduce a method for preserving colors in non-editing regions, which are automatically detected by the model. Together, Sec. 3.3 and Sec. 3.4 constitute the core of our method. Finally, Sec. 3.5 introduces word-level control over color attributes via attention re-weighting, serving as an optional add-on for fine-grained attribute control.
@@ 3.1. Task Formulation: Text-guided Color Editing
  • $G=\{G_{1},\dots,G_{N}\}$ - a set of object geometries, e.g., mesh topology for NN objects,
@@
    Material consistency - the object material $S$ remain fixed. The edits apply only to the albedo components AA of specific objects.
@@
Figure 2. Pipeline of ColorCtrl. (a) Visualizes the attention mechanism in $M$-DiT blocks. (b) Enables effective color editing while maintaining structural consistency. (c) Preserves colors in non-editing regions. (d) Applies attribute re-weighting to specific word tokens. Symbols in the source branch have no superscript. Symbols with a superscript indicate the target branch, and hats (e.g., $\hat{V}$, $\hat{M}$) denote edited outputs.
3.2. Preliminaries: Attention in $M$-DiT Blocks
The attention mechanism for text-vision fusion in $M$-DiT (Esser et al., 2024) differs from that in U-Net (Rombach et al., 2022). In U-Net, cross-attention is used for text guidance, while self-attention focuses on visual content interactions, separately. In contrast, $M$-DiT integrates text and vision by concatenating their tokens together and processing them jointly via sole self-attention. As illustrated in Fig. 2 (a), vision and text tokens $\bm{z}$ are fed into the $i$-th $M$-DiT block at timestep $t$ during inference. After modulation, the block produces an attention map $\bm{M}$ and value tokens $\bm{V}$, which are used to generate the updated tokens $\hat{z}$ for the next block and timestep. The resulting attention map $\bm{M}$ can be divided into four parts: vision-to-vision (upper-left), vision-to-text (upper-right), text-to-vision (lower-left), and text-to-text (lower-right). These quadrants correspond to different query-key token pairings. For example, the text-to-vision region represents attention scores computed between query tokens from the text modality $\bm{Q^{\text{text}}}$ and key tokens from the vision modality $\bm{K^{\text{vision}}}$. Similarly, the value tokens $\bm{V}$ are composed of a vision part $\bm{V^{\text{vision}}}$ and a text part $\bm{V^{\text{text}}}$. The functional roles of each region are discussed in detail in the following sections.
$3.$ Structure Preservation
@@ 3.5. Attribute Re-Weighting
So far, our method enables powerful color editing capabilities. However, the granularity of control via text prompts remains limited. For example, if the user specifies a color like “dark yellow”, there is no way to explicitly control the degree of darkness. Existing re-weighting techniques in U-Net-based models primarily follow two approaches: (1) Scaling the text embedding of specific tokens before the diffusion process (Perry, 2023). However, this method is designed for CLIP-based (Radford et al., 2021) text encoders and is incompatible with $M$-DiT-based models, which typically use T5 (Raffel et al., 2020). (2) Scaling the attention scores of specific tokens in the cross-attention map (Hertz et al., 2023). This is also inapplicable to $M$-DiT, which relies solely on self-attention. Moreover, both approaches fail to maintain structural consistency during scaling, limiting their utility in high-fidelity editing tasks.
To support more fine-grained and user-friendly control in $M$-DiT-based models, we introduce a mechanism to modulate the strength of specific attribute words. Specifically, we scale the attention scores in the text-to-vision parts of the attention map corresponding to the selected word tokens before the softmax operation, as illustrated in Fig. 2 (d). This modification can be applied either directly on the original source attention maps or on the target attention maps that have already undergone structure preservation, allowing flexible integration in above editing pipeline.
@@ 4.1. Setup
Baselines. We compare our method against several training-free approaches built upon $M$-DiT, including FireFlow (Deng et al., 2025), RF-Solver (Wang et al., 2025), SDEdit (Meng et al., 2022), DiTCtrl (Cai et al., 2025), FlowEdit (Kulikov et al., 2024), and UniEdit-Flow (Jiao et al., 2025). We focus exclusively on $M$-DiT-based baselines, as prior work (Jiao et al., 2025; Deng et al., 2025) has shown that U-Net-based methods perform significantly worse. Accordingly, we exclude methods that cannot be adapted to the $M$-DiT architecture. For commercial baselines, we compare with FLUX.1 Kontext Max (Black Forest Labs, 2025) and GPT-4o Image Generation (OpenAI, 2025).
@@ 4.3. Comparison with Commercial Models (Images)
$4.$ Comparison with Training-Free Methods (Videos)
@@ 5. Conclusion
We have introduced ColorCtrl, a training-free method for text-guided color editing that achieves fine-grained, physically consistent control over albedo, light source color, and ambient illumination. Our method is designed to edit only the intended visual attributes specified in the prompt, leaving all unrelated regions untouched. Built upon diverse $M$-DiT-based diffusion models, such as SD3 and FLUX.1-dev, our approach enables fine-grained control over attribute intensity, while preserving geometry, material properties, and light-matter interactions. ColorCtrl not only outperforms prior training-free methods and achieves state-of-the-art results but also delivers stronger consistency than commercial models, i.e., FLUX.1 Kontext Max and GPT-4o Image Generation, on both quantitative and qualitative evaluations. Moreover, its model-agnostic design generalizes naturally to video diffusion models (i.e., CogVideoX) and instruction-based editing diffusion models (i.e., Step1X-Edit and FLUX.1 Kontext dev), highlighting its broad applicability. We believe ColorCtrl paves the way for scalable, high-fidelity, and controllable color editing in both research and practical deployment.
@@ A.3. Implementation of Compared Methods
  • DiTCtrl (Cai et al., 2025): For SD3-based image editing, we set the editing range from timestep 22 to $17$, modifying the last 55 blocks. For FLUX.1-dev, we edit from timestep 22 to $11$ across the last 66 blocks. For video editing with CogVideoX-2B, the official implementation is used. During editing, key ($\bm{K}$)and value ($\bm{V}$) tokens from the source branch are copied into the attention layers of the target branch.
@@
  • UniEdit-Flow (Jiao et al., 2025): The official version supports SD3 and FLUX.1-dev but provides the $\omega$ parameter only for CFG = 11. Following the similarity transformation introduced in the paper, we use $\omega=5\div 7.5\approx 0.6$ and set $\alpha=0.6$ for SD3, $\alpha=0.85$ for FLUX.1-dev, and $\alpha=0.8$ for video generation, which yields comparable performance to the original.
  • FireFlow (Deng et al., 2025): In SD3, we observe that it is difficult to select a suitable end timestep for FireFlow, as setting it too high often leads to artifacts and generation failures, as shown in Fig. 9. To mitigate quality degradation caused by excessive editing steps, we restrict editing to timesteps 0 through 3 across all blocks in SD3. In contrast, FLUX.1-dev does not exhibit this issue, and using the same range (timesteps 0 to 3) yields results comparable to those reported in the original paper. For video generation, the editing ends at timestep 99. During editing, value tokens ($\bm{V}$) are copied from the source to the target.
  • RF-Solver (Wang et al., 2025): RF-Solver exhibits a similar issue to FireFlow when applied to SD3, where higher editing timesteps lead to artifacts and degraded generation quality, as illustrated in Fig. 9. To address this, we limit editing to timesteps 0 to 7 in the second half of the SD3 blocks. For FLUX.1-dev, we set the end timestep to 4, which yields stable results without noticeable artifacts. Value tokens ($\bm{V}$) are copied from the source to the target. For video generation, editing is performed up to timestep 99.

### v2

@@ Abstract.
Text-guided color editing in images and videos is a fundamental yet unsolved problem, requiring fine-grained manipulation of color attributes, including albedo, light source color, and ambient lighting, while preserving physical consistency in geometry, material properties, and light-matter interactions. Existing training-free methods offer broad applicability across editing tasks but struggle with precise color control and often introduce visual inconsistency in both edited and non-edited regions. In this work, we present ColorCtrl, a training-free color editing method that leverages the attention mechanisms of modern Multi-Modal Diffusion Transformers (MM-DiT). By disentangling structure and color through targeted manipulation of attention maps and value tokens, our method enables accurate and consistent color editing, along with word-level control of attribute intensity. Our method modifies only the intended regions specified by the prompt, leaving unrelated areas untouched. Extensive experiments on both SD3 and FLUX.1-dev demonstrate that ColorCtrl outperforms existing training-free approaches and achieves state-of-the-art performances in both edit quality and consistency. Furthermore, our method surpasses strong commercial models such as FLUX.1 Kontext Max and GPT-4o Image Generation in terms of consistency. When extended to video models like CogVideoX, our approach exhibits greater advantages, particularly in maintaining temporal coherence and editing stability. Finally, our method also generalizes to instruction-based editing diffusion models such as Step1X-Edit and FLUX.1 Kontext dev, further demonstrating its versatility. Here is the project website.
copyright: nonecopyright: none
@@ 1. Introduction
The recent architectural shift in diffusion models from U-Net (Rombach et al., 2022) to the Multi-Modal Diffusion Transformer (MM-DiT) (Esser et al., 2024) opens up new opportunities for training-free color editing, for two main reasons: (1) the new architecture allows for scaling up both data and model capacity, enabling better adherence to physical priors; and (2) the improved fusion of text and vision modalities supports more flexible and precise attention-control editing strategies.
In this work, we propose ColorCtrl, a training-free, open-world color editing method that effectively leverages pre-trained MM-DiT models to perform natural and precise color modifications, while preserving all other visual attributes, such as geometry, material, reflection and refraction, light source position, illumination direction, and light intensity, as shown in Fig. 1. For example, when editing the color of a ball, ColorCtrl accurately adjusts not only the ball itself but also its reflection in the water, the specular reflections on both sides, and even the small water droplets on the glass surface, as demonstrated in the fourth example of Fig. 1. Furthermore, ColorCtrl supports fine-grained, word-level control over the strength of color attributes while preserving the other visual attributes mentioned above. Thanks to our precise and robust control, ColorCtrl requires no manual parameter tuning and can be directly applied across all attention layers and inference timesteps.
@@
  • We present extensive experiments demonstrating that ColorCtrl achieves state-of-the-art performance among training-free methods on MM-DiT-based models. Compared to commercial models (i.e., FLUX.1 Kontext Max and GPT-4o Image Generation), our method delivers significantly better consistency preservation and produces more natural edits.
  • ColorCtrl generalizes well across multiple MM-DiT-based models, including video and instruction-based editing models, highlighting its broad applicability and extensibility.
@@ 2. Related Work
Text-to-image and video generation. Diffusion models with a U-Net backbone (Ho et al., 2020; Rombach et al., 2022; Guo et al., 2024) have largely replaced early GAN systems (Reed et al., 2016; Yu et al., 2023; Wang et al., 2023) due to superior image fidelity. However, the U-Net design scales poorly, leading to the adoption of Diffusion Transformers (DiT) (Peebles and Xie, 2023). Among these, MM-DiT (Esser et al., 2024) has emerged as a widely adopted backbone for recent state-of-the-art models (Esser et al., 2024; AI, 2024; Labs, 2024; Yang et al., 2024; Kong et al., 2024; Liu et al., 2025b; Black Forest Labs, 2025), such as SD3 (Esser et al., 2024) and FLUX.1-dev (Labs, 2024) for image generation, as well as CogVideoX (Yang et al., 2024) for video generation. In this work, we propose an attention-control scheme that plugs into any MM-DiT model.
Text-guided editing. Training-free text-guided editing methods based on pre-trained diffusion models offer flexibility and efficiency. Existing approaches fall into two groups: (i) sampling-based methods, which steer generation by injecting controlled noise or inversion (Jiao et al., 2025; Huberman-Spiegelglas et al., 2024; Xu et al., 2023; Kulikov et al., 2024; Yan et al., 2025); (ii) attention-based methods, which modify attention maps, starting with Prompt-to-Prompt (Hertz et al., 2023) and its image/video variants (Chen et al., 2024; Wang et al., 2025; Liu et al., 2024; Cao et al., 2023; Cai et al., 2025; Rout et al., 2025; Xu et al., 2025). Among these, DiTCtrl is the only method exploring attention control in MM-DiT, but ColorCtrl differs in key ways: DiTCtrl only applies mask extraction during long video generation, not color editing, while our improved method is more robust. Additionally, DiTCtrl’s attribute re-weighting disrupts attention consistency, causing inconsistent geometries, while our method avoids this. Furthermore, ColorCtrl works across all layers and timesteps, unlike DiTCtrl, which requires careful layer selection. Different from DiTCtrl, we operate in attention maps rather than key and value tokens. More recently, methods (Liu et al., 2025a; Brooks et al., 2023) like FLUX.1 Kontext (Black Forest Labs, 2025) and GPT-4o (OpenAI, 2025) have shown convenient creative workflows in training with synthetic instruction-response pairs for fine-tuning a diffusion model for image editing.
@@ 3. Method
We first formulate the color editing problem in Sec. 3.1, followed by a revisit of the attention mechanism in MM-DiT blocks in Sec. 3.2. Sec. 3.3 describes our approach for preserving geometry, material properties, and light-matter interactions. In Sec. 3.4, we introduce a method for preserving colors in non-editing regions, which are automatically detected by the model. Together, Sec. 3.3 and Sec. 3.4 constitute the core of our method. Finally, Sec. 3.5 introduces word-level control over color attributes via attention re-weighting, serving as an optional add-on for fine-grained attribute control.
@@ 3.1. Task Formulation: Text-guided Color Editing
  • $G=\{G_{1},\dots,G_{N}\}$ - a set of object geometries, e.g., mesh topology for $N$ objects,
@@
    Material consistency - the object material $S$ remain fixed. The edits apply only to the albedo components $A$ of specific objects.
@@
Figure 2. Pipeline of ColorCtrl. (a) Visualizes the attention mechanism in MM-DiT blocks. (b) Enables effective color editing while maintaining structural consistency. (c) Preserves colors in non-editing regions. (d) Applies attribute re-weighting to specific word tokens. Symbols in the source branch have no superscript. Symbols with a superscript indicate the target branch, and hats (e.g., $\hat{V}$, $\hat{M}$) denote edited outputs.
3.2. Preliminaries: Attention in MM-DiT Blocks
The attention mechanism for text-vision fusion in MM-DiT (Esser et al., 2024) differs from that in U-Net (Rombach et al., 2022). In U-Net, cross-attention is used for text guidance, while self-attention focuses on visual content interactions, separately. In contrast, MM-DiT integrates text and vision by concatenating their tokens together and processing them jointly via sole self-attention. As illustrated in Fig. 2 (a), vision and text tokens $\bm{z}$ are fed into the $i$-th MM-DiT block at timestep $t$ during inference. After modulation, the block produces an attention map $\bm{M}$ and value tokens $\bm{V}$, which are used to generate the updated tokens $\hat{z}$ for the next block and timestep. The resulting attention map $\bm{M}$ can be divided into four parts: vision-to-vision (upper-left), vision-to-text (upper-right), text-to-vision (lower-left), and text-to-text (lower-right). These quadrants correspond to different query-key token pairings. For example, the text-to-vision region represents attention scores computed between query tokens from the text modality $\bm{Q^{\text{text}}}$ and key tokens from the vision modality $\bm{K^{\text{vision}}}$. Similarly, the value tokens $\bm{V}$ are composed of a vision part $\bm{V^{\text{vision}}}$ and a text part $\bm{V^{\text{text}}}$. The functional roles of each region are discussed in detail in the following sections.
3.3. Structure Preservation
@@ 3.5. Attribute Re-Weighting
So far, our method enables powerful color editing capabilities. However, the granularity of control via text prompts remains limited. For example, if the user specifies a color like “dark yellow”, there is no way to explicitly control the degree of darkness. Existing re-weighting techniques in U-Net-based models primarily follow two approaches: (1) Scaling the text embedding of specific tokens before the diffusion process (Perry, 2023). However, this method is designed for CLIP-based (Radford et al., 2021) text encoders and is incompatible with MM-DiT-based models, which typically use T5 (Raffel et al., 2020). (2) Scaling the attention scores of specific tokens in the cross-attention map (Hertz et al., 2023). This is also inapplicable to MM-DiT, which relies solely on self-attention. Moreover, both approaches fail to maintain structural consistency during scaling, limiting their utility in high-fidelity editing tasks.
To support more fine-grained and user-friendly control in MM-DiT-based models, we introduce a mechanism to modulate the strength of specific attribute words. Specifically, we scale the attention scores in the text-to-vision parts of the attention map corresponding to the selected word tokens before the softmax operation, as illustrated in Fig. 2 (d). This modification can be applied either directly on the original source attention maps or on the target attention maps that have already undergone structure preservation, allowing flexible integration in above editing pipeline.
@@ 4.1. Setup
Baselines. We compare our method against several training-free approaches built upon MM-DiT, including FireFlow (Deng et al., 2025), RF-Solver (Wang et al., 2025), SDEdit (Meng et al., 2022), DiTCtrl (Cai et al., 2025), FlowEdit (Kulikov et al., 2024), and UniEdit-Flow (Jiao et al., 2025). We focus exclusively on MM-DiT-based baselines, as prior work (Jiao et al., 2025; Deng et al., 2025) has shown that U-Net-based methods perform significantly worse. Accordingly, we exclude methods that cannot be adapted to the MM-DiT architecture. For commercial baselines, we compare with FLUX.1 Kontext Max (Black Forest Labs, 2025) and GPT-4o Image Generation (OpenAI, 2025).
@@ 4.3. Comparison with Commercial Models (Images)
4.4. Comparison with Training-Free Methods (Videos)
@@ 5. Conclusion
We have introduced ColorCtrl, a training-free method for text-guided color editing that achieves fine-grained, physically consistent control over albedo, light source color, and ambient illumination. Our method is designed to edit only the intended visual attributes specified in the prompt, leaving all unrelated regions untouched. Built upon diverse MM-DiT-based diffusion models, such as SD3 and FLUX.1-dev, our approach enables fine-grained control over attribute intensity, while preserving geometry, material properties, and light-matter interactions. ColorCtrl not only outperforms prior training-free methods and achieves state-of-the-art results but also delivers stronger consistency than commercial models, i.e., FLUX.1 Kontext Max and GPT-4o Image Generation, on both quantitative and qualitative evaluations. Moreover, its model-agnostic design generalizes naturally to video diffusion models (i.e., CogVideoX) and instruction-based editing diffusion models (i.e., Step1X-Edit and FLUX.1 Kontext dev), highlighting its broad applicability. We believe ColorCtrl paves the way for scalable, high-fidelity, and controllable color editing in both research and practical deployment.
@@ A.3. Implementation of Compared Methods
  • DiTCtrl (Cai et al., 2025): For SD3-based image editing, we set the editing range from timestep $2$ to $17$, modifying the last $5$ blocks. For FLUX.1-dev, we edit from timestep $2$ to $11$ across the last $6$ blocks. For video editing with CogVideoX-2B, the official implementation is used. During editing, key ($\bm{K}$)and value ($\bm{V}$) tokens from the source branch are copied into the attention layers of the target branch.
@@
  • UniEdit-Flow (Jiao et al., 2025): The official version supports SD3 and FLUX.1-dev but provides the $\omega$ parameter only for CFG = $1$. Following the similarity transformation introduced in the paper, we use $\omega=5\div 7.5\approx 0.6$ and set $\alpha=0.6$ for SD3, $\alpha=0.85$ for FLUX.1-dev, and $\alpha=0.8$ for video generation, which yields comparable performance to the original.
  • FireFlow (Deng et al., 2025): In SD3, we observe that it is difficult to select a suitable end timestep for FireFlow, as setting it too high often leads to artifacts and generation failures, as shown in Fig. 9. To mitigate quality degradation caused by excessive editing steps, we restrict editing to timesteps 0 through 3 across all blocks in SD3. In contrast, FLUX.1-dev does not exhibit this issue, and using the same range (timesteps 0 to 3) yields results comparable to those reported in the original paper. For video generation, the editing ends at timestep $9$. During editing, value tokens ($\bm{V}$) are copied from the source to the target.
  • RF-Solver (Wang et al., 2025): RF-Solver exhibits a similar issue to FireFlow when applied to SD3, where higher editing timesteps lead to artifacts and degraded generation quality, as illustrated in Fig. 9. To address this, we limit editing to timesteps 0 to 7 in the second half of the SD3 blocks. For FLUX.1-dev, we set the end timestep to 4, which yields stable results without noticeable artifacts. Value tokens ($\bm{V}$) are copied from the source to the target. For video generation, editing is performed up to timestep $9$.